Solve the following equations

2sin2x-3cosx = 0

sinx + cosx =0



1
Expert's answer
2021-09-02T09:07:39-0400

"2sin2x-3cosx\:=\:0\:"

"-3\cos \:\left(x\right)+4\cos \left(x\right)\sin \left(x\right)=0"

"\cos \left(x\right)\left(4\sin \left(x\right)-3\right)=0"

"\cos \left(x\right)=0\quad \:\mathrm{or}\quad \:\:4\sin \left(x\right)-3=0"

"x=\frac{\pi \:}{2}+2\pi \:n,\:x=\frac{3\pi \:}{2}+2\pi \:n,\:x=\arcsin \:\left(\frac{3}{4}\right)+2\pi \:n,\:x=\pi \:-\arcsin \:\left(\frac{3}{4}\right)+2\pi \:n"

"x=\frac{\pi }{2}+2\pi n,\:x=\frac{3\pi }{2}+2\pi n,\:x=0.84806\dots +2\pi n,\:x=\pi -0.84806\dots +2\pi n"


"sinx\:+\:cosx\:=0"

"\tan \left(x\right)+1=0"

"\tan \left(x\right)+1-1=0-1"

"\tan \left(x\right)=-1"

"x=\frac{3\pi \:}{4}+\pi \:n"